An​ on-line retailer identified the web browser being used by a sample of 50 shoppers to its online site. The accompanying data table identifies the browser being used by a shopper.

Previously in​ 2010, 64​% of shoppers used Browser​ A, 25​% Browser​ B, 5​% Browser​ C, 3​% Browser​ D, and 3​% Browser E. Complete parts​ (a) through​ (c) below.

(a) Using​ software, tabulate the frequency of the choice of browser used by these shoppers.

Browser A= ? shoppers
Answer #1

(A) total number of shoppers (n) = 50

proportion of Browser A shoppers = 25/100 = 0.25

Frequency for Browser A = n*p

= 50*0.25

= 12.5